You may need to draw the figures indicated on your scratch paper to answer this question.
Find the slope of the line P (- 6, 8), Q (- 6, 12).
m = ?

Answer by stanbon(75887)

Find the slope of the line P (- 6, 8), Q (- 6, 12).
---
slope = change in y/change in x
---

m = (12-8)/(-6--6) = 4/0
The slope is "undefined".
---------------
Cheers,
Stan H.